1. find the equation of the line passing through points (6,-1) and (5,2). write your answer in slope intercept form
2. find the equation of the line passing through (-3,10) that is perpendicular to the line y=(-3/2)x+4 and graph the line

Answer by stanbon(75887)

1. find the equation of the line passing through points (6,-1) and (5,2). write your answer in slope intercept form
------
slope = (2--1)/(5-6) = -3/1 = -3
intercept: 2 = -3*5 + b
b = 17
Equation:
y = -3 -17
-------------------------------------------
2. find the equation of the line passing through (-3,10) that is perpendicular to the line y=(-3/2)x+4 and graph the line
The given line has slope = -3/2
Any perpendicular line must have slope = 2/3
---
intercept : 10 = (2/3)(-3) = b
b = 12
Equation:
y = (2/3)x + 12
========================
Cheers,
Stan H.